"Suppose that one cocktail is 25% ethanol and another cocktail is 110% ethanol. How many milliliters of each solution should be mixed to make 10.5 milliliters of an 80% ethanol solution"?
Can I get help with setting up an equation and solving please?
Thank you for your time.

Answer by rothauserc(4718)

we are given
.25x solution and a 1.10y solution
x+y = 10.5 mL
.80 * 10.5 = 8.4 mL
now we can solve for x
x = 10.5 - y and substituting for x we have
.25*(10.5-y) + 1.10*y = 8.4
2.65 -.25y + 1.10y = 8.4
.85y = 5.75
y = 6.76 mL and
x = 10.5 - 6.76 = 3.74 mL
check the answers
.25(3.74) + 1.10(6.76) = 8.4
.935 + 7.436 = 8.4
8.4 = 8.4